### v2

# Partial-Label Learning with a Reject Option

Abstract
In real-world applications, one often encounters ambiguously labeled data, where different annotators assign conflicting class labels. Partial-label learning allows training classifiers in this weakly supervised setting, where state-of-the-art methods already show good predictive performance. However, even the best algorithms give incorrect predictions, which can have severe consequences when they impact actions or decisions. We propose a novel risk-consistent partial-label learning algorithm with a reject option, that is, the algorithm can reject unsure predictions. Extensive experiments on artificial and real-world datasets show that our method provides the best trade-off between the number and accuracy of non-rejected predictions when compared to our competitors, which use confidence thresholds for rejecting unsure predictions instead. When evaluated without the reject option, our nearest neighbor-based approach also achieves competitive prediction performance.
1 Introduction
Real-world data is often noisy, for example, human annotators might assign different class labels to the same instance. In partial-label learning (PLL; Hüllermeier & Beringer 2005; Liu & Dietterich 2012; Zhang & Yu 2015; Xu et al. 2023), training instances can have multiple labels, known as candidates, of which only one is correct. While in some cases it is possible to sanitize such data, cleaning is costly, especially for large-scale datasets. Instead, one wants to predict the class labels of unseen instances having sets of candidates only, that is, without knowing the exact ground-truth labels of the training data. PLL algorithms allow for handling such ambiguously labeled data.
However, even the best algorithms give incorrect predictions. These errors can have severe consequences when they impact actions or decisions. Consider, for example, safety-critical domains such as the classification of medical images (Yang et al., 2009; Lambrou et al., 2011; Senge et al., 2014; Kendall & Gal, 2017; Reamaroon et al., 2019) or the control of self-driving cars (Xu et al., 2014; Varshney & Alemzadeh, 2017; Hubmann et al., 2017; Shafaei et al., 2018; Michelmore et al., 2020). One option to limit fallacies is to refuse to give an individual prediction if unsure and, instead, let humans decide on the label of an instance or the actions to take. A reject option allows one to abstain from certain predictions based on a threshold. Naturally, there is a trade-off arising between the number and accuracy of non-rejected predictions. In the supervised setting, reject options have already been studied, both, for classification (Hellman, 1970; Herbei & Wegkamp, 2006) and regression tasks (Zaoui et al., 2020). In PLL, a reject option can also mitigate misclassifications, which are more likely due to the weak supervision. We are the first to study reject options in PLL.
Determining whether to reject a prediction of a PLL algorithm is highly non-trivial as the uncertainty involved in prediction-making originates from multiple sources (Hora, 1996). On the one hand, (i) there is inherent uncertainty due to ambiguously labeled data (aleatoric uncertainty). This ambiguity is influenced by both instances and classes: In some regions of the instance space, annotators are more likely to mislabel instances than in others. Additionally, some class labels are inherently more similar than others. On the other hand, (ii) there is uncertainty due to the lack of knowledge regarding the relevance of each candidate label (epistemic uncertainty). Optimally, an algorithm does not mix (i) and (ii) in deciding whether to reject a prediction.
In fact, Fisher (1930, 1935) already notes that “not knowing the chance of mutually exclusive events [epistemic uncertainty] and knowing the chance to be equal [aleatoric uncertainty] are two quite different states of knowledge.” One can reduce epistemic uncertainty with more data or better models; aleatoric uncertainty is irreducible (Meyerowitz et al., 1994; Harmanec et al., 1996; Klir & Wierman, 1999; Abellán et al., 2006; Sale et al., 2023; Wimmer et al., 2023). To our knowledge, existing partial-label learning approaches, for example, (Hüllermeier & Beringer, 2005; Liu & Dietterich, 2012; Zhang & Yu, 2015; Ni et al., 2021; Xu et al., 2023), model label information using standard probability, which does not allow to discern aleatoric and epistemic uncertainty.
We propose a partial-label learning approach based on Dempster-Shafer theory that keeps both uncertainties separate and achieves a significantly better trade-off of the number and accuracy of non-rejected predictions when compared to the state-of-the-art.
Contributions. We summarize our contributions as follows.
  • We introduce Dst-Pll, a novel partial-label learning algorithm with a reject option that learns from ambiguously labeled data. The algorithm effectively determines whether to reject a prediction and provides the best trade-off between the number and accuracy of non-rejected predictions.
  • Extensive experiments on artificial and real-world data support our claims. We make our code and data openly available. [^1] https://github.com/anon1248/pll-with-a-reject-option.
  • We analyze Dst-Pll, give a closed-form expression of its expected decision boundary, and prove its risk consistency. We believe these results are of independent interest. The runtime analysis shows that the proposed method is dominated by $k$ nearest-neighbor search, which has an average time complexity of $\mathcal{O}(dk\log n)$ with $d$ features and $n$ training instances.
Structure of the paper. We define the problem setting in Section 2 and Section 3 discusses related work. We propose our method in Section 4, Section 5 features experiments and Section 6 concludes. All proofs and additional experiments are in the appendices.
2 Problem Statement and Notations
This section formally defines the partial-label learning problem, establishes notations used throughout this work, introduces reject options, and briefly summarizes Dempster-Shafer theory.
Partial-Label Learning (PLL) and Notations. Let $\operatorname{\mathcal{X}}=\mathbb{R}^{d}$ denote a $d$-dimensional real-valued feature space and $\operatorname{\mathcal{Y}}=[l]:=\{1,\dots,l\}$ the finite set of $3\leq l\in\mathbb{N}$ classes. A partial-label learning training dataset $\mathcal{D}=\{(x_{i},s_{i})\mid i\in[n]\}$ of $n$ instances contains feature vectors $x_{i}\in\operatorname{\mathcal{X}}$ and a set of candidate labels $s_{i}\subseteq\operatorname{\mathcal{Y}}$ for each $i\in[n]$. All instances $i$ have an unknown ground-truth label $y_{i}\in\operatorname{\mathcal{Y}}$, and $y_{i}\in s_{i}$. Further, the candidate labels $s_{i}$ can be partitioned into $s_{i}=\{y_{i}\}\cup z_{i}$ with $y_{i}\notin z_{i}$, that is, $z_{i}\subset\mathcal{Y}$ are the false-positive labels.
The sample space is $\operatorname{\Omega}=\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\times 2^{\operatorname{\mathcal{Y}}}$. It is equipped with the smallest Borel $\sigma$-algebra $\operatorname{\mathcal{B}}(\operatorname{\Omega})$, yielding the measurable space $(\operatorname{\Omega},\operatorname{\mathcal{B}}(\operatorname{\Omega}))$ with respect to which the set of all probability measures $\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\Omega},\operatorname{\mathcal{B}}(\operatorname{\Omega}))$ and a probability measure $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\Omega},\operatorname{\mathcal{B}}(\operatorname{\Omega}))$ is defined. The random variables $X:\operatorname{\Omega}\rightarrow\operatorname{\mathcal{X}}$, $Y:\operatorname{\Omega}\rightarrow\operatorname{\mathcal{Y}}$, and $S:\operatorname{\Omega}\rightarrow 2^{\operatorname{\mathcal{Y}}}$ govern an instance’s features, true label, and candidate labels, respectively. We denote the corresponding marginal measures by associating the corresponding index to $\operatorname{\mathbb{P}}$, for example, $\operatorname{\mathbb{P}}_{X}$ is the distribution of $X$. $\operatorname{\mathbb{P}}_{XY}$ denotes the joint distribution of $(X,Y)$.
PLL aims to train a classifier $g:\operatorname{\mathcal{X}}\to\operatorname{\mathcal{Y}}$ that, given a loss function $\operatorname{\mathcal{L}}:\operatorname{\mathcal{Y}}\times 2^{\operatorname{\mathcal{Y}}}\to\mathbb{R}_{\geq 0}$, minimizes the risk $\operatorname{\mathcal{R}}(g)=\operatorname{\mathbb{E}}_{(X,S)\sim\operatorname{\mathbb{P}}_{XS}}\operatorname{\mathcal{L}}(g(X),S)$. Common choices for $\operatorname{\mathcal{L}}$ include the average loss (Nguyen & Caruana, 2008; Cour et al., 2011) and the minimum loss (Lv et al., 2020; Xu et al., 2021). The empirical version of the risk substitutes the expectation with a sample mean.
Reject Option. A reject option $\Gamma_{g}:\operatorname{\mathcal{X}}\to 2^{\operatorname{\mathcal{Y}}}$ associated with a trained classifier $g$ either returns $g$’s prediction (accept) or abstains from making any prediction at all (reject), that is, $\Gamma_{g}(x)\in\{\emptyset,\{g(x)\}\}$ for $x\in\operatorname{\mathcal{X}}$ with $\Gamma_{g}(x)=\emptyset$ denoting a reject. We then define the rejection rate $\operatorname{r}(\Gamma_{g})=\operatorname{\mathbb{P}}_{X}(\Gamma_{g}(X)=\emptyset)$ and the expected error $\operatorname{err}(\Gamma_{g})=\operatorname{\mathbb{E}}_{(X,S)\sim\operatorname{\mathbb{P}}_{XS}}[\operatorname{\mathcal{L}}(g(X),S)\mathds{1}_{\{\Gamma_{g}(X)\neq\emptyset\}}]$ of accepted predictions. Naturally, a trade-off arises between the number and accuracy of accepted predictions, which leads us to the risk
$\operatorname{\mathcal{R}}_{\lambda}(\Gamma_{g})=\operatorname{\mathbb{E}}_{(X,S)\sim\operatorname{\mathbb{P}}_{XS}}[\operatorname{\mathcal{L}}(g(X),S)\mathds{1}_{\{\Gamma_{g}(X)\neq\emptyset\}}+\lambda\mathds{1}_{\{\Gamma_{g}(X)=\emptyset\}}]=\operatorname{err}(\Gamma_{g})+\lambda\operatorname{r}(\Gamma_{g})\text{,}$ (1)
for $\lambda\geq 0$. $\operatorname{\mathcal{L}}$ characterizes the cost of misclassification and parameter $\lambda$ the cost of rejecting a prediction. This setting has already been considered in supervised classification (Hellman, 1970; Herbei & Wegkamp, 2006) and in regression (Zaoui et al., 2020). We are the first to study it in PLL.
[…]
$\operatorname{bel}_{\operatorname{m}}(A):=\sum_{B\subseteq A}\operatorname{m}(B)\quad\text{and}\quad\operatorname{pl}_{\operatorname{m}}(A):=\sum_{B\subseteq\operatorname{\mathcal{Y}},A\cap B\neq\emptyset}\operatorname{m}(B)\quad\text{for}\ A\in 2^{\operatorname{\mathcal{Y}}}\text{,}$ (2)
the set of all probability measures supporting $\operatorname{m}$ is
$\operatorname{\mathcal{C}}_{\operatorname{m}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}}):=\big{\{}\operatorname{\mathbb{P}}\in\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})\mid\operatorname{bel}_{\operatorname{m}}(A)\leq\operatorname{\mathbb{P}}(A)\leq\operatorname{pl}_{\operatorname{m}}(A)\ \text{for all $A\subseteq\operatorname{\mathcal{Y}}$}\big{\}}\text{.}$ (3)
The set $\operatorname{\mathcal{C}}_{\operatorname{m}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})\subseteq\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ is called a credal set (Abellán et al., 2006; Cuzzolin, 2021). Further, DST provides rules to combine $\operatorname{m}$-s from multiple sources (Dempster, 1967; Yager, 1987a, b).
3 Related Work
Imperfect data often renders the application of supervised methods challenging. Weakly supervised learning tackles this setting and encompasses a variety of problem formulations (Bylander, 1994; Hady & Schwenker, 2013; Ishida et al., 2019) including PLL. We discuss related work regarding PLL in Section 3.1. Section 3.2 elaborates on related work in dealing with uncertainty in ML.
3.1 Partial-Label Learning
The early PLL approaches transfer common supervised learning frameworks to the PLL context: Grandvalet (2002) proposes a logistic regression formulation, Jin & Ghahramani (2002) propose an expectation-maximization strategy, Hüllermeier & Beringer (2005) propose a $k$-nearest-neighbors method, Nguyen & Caruana (2008) propose an extension of support-vector machines, and Cour et al. (2011) introduce an average loss formulation allowing for the use of any supervised method. However, these approaches (i) cannot model the relevance of candidate labels in the labeling process or (ii) are not robust to non-uniform noise in the candidate sets. We re-emphasize that (i) and (ii) stem from two different sources of uncertainty: (i) arises from the lack of knowledge about the label relevancies of the PLL model (epistemic uncertainty) and (ii) arises from the random label noise in the candidate sets and is inherent to the PLL problem (aleatoric uncertainty).
More recent approaches address (i) and (ii) in the following ways: Zhang & Yu (2015); Zhang et al. (2016); Xu et al. (2019); Wang et al. (2019); Feng & An (2019); Ni et al. (2021) leverage ideas from representation learning (Bengio et al., 2013), Yu & Zhang (2017); Wang et al. (2019); Feng & An (2019); Ni et al. (2021) extend the maximum-margin idea, Liu & Dietterich (2012); Lv et al. (2020) propose extensions of the expectation-maximization strategy, Zhang et al. (2017); Tang & Zhang (2017); Wu & Zhang (2018) propose stacking and boosting ensembles, Lv et al. (2020); Cabannes et al. (2020) introduce a minimum loss formulation, and Feng et al. (2020); Lv et al. (2020); Xu et al. (2021); Wang et al. (2022); He et al. (2022); Xu et al. (2023); Fan et al. (2024) employ deep-learning techniques such as auto-encoders, data augmentation, and transformers in the context of PLL. These methods address (i) and (ii), featuring already good performance. However, considering a reject option of those methods based on a confidence threshold, as proposed in the supervised setting (Herbei & Wegkamp, 2006), yields sub-optimal results as all of the listed methods mix epistemic and aleatoric uncertainty stemming from (i) and (ii), respectively. We claim that keeping both separate provides a better trade-off for the number and accuracy of non-rejected predictions (Section 5).
3.2 Dealing with Uncertainty
Recently, much attention has been given to quantifying the uncertainty associated with predictions by machine-learning models. These include Bayesian methods (Kingma & Welling, 2014; Kendall & Gal, 2017) and ensembles (Lakshminarayanan et al., 2017; Wimmer et al., 2023). Notably, a probability measure $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ cannot distinguish uncertainty arising from (i) and (ii) (Hüllermeier & Waegeman, 2021). To mitigate this limitation, we employ Dempster-Shafer theory (Dempster, 1967; Shafer, 1986). We argue that DST is a natural fit for PLL: With DST, we can assign probability mass to sets of candidates without specifying the probabilities of individual labels; incorrect labels do not obtain any probability mass. In contrast, existing PLL approaches (Hüllermeier & Beringer, 2005; Cour et al., 2011; Liu & Dietterich, 2012; Zhang & Yu, 2015; Ni et al., 2021; Xu et al., 2023) initially attribute some probability to each label candidate and subsequently refine them. By doing so, most probability mass is initially allocated to labels that are certainly incorrect, as only one candidate is the true label. Several methods already leverage DST in supervised learning (Mandler & Schümann, 1988; Denoeux, 1995; Tabassian et al., 2012; Sensoy et al., 2018; Denoeux, 2019; Tong et al., 2021). We consider the nearest neighbor approach by Denoeux (1995) to be most closely related to our approach. Basic probability assignments are constructed from the nearest neighbors of an instance. Then, Dempster’s rule is used to combine them into a single bpa. Their analysis is, however, not transferable to our case because they only have singletons or the full label space as focal sets making set intersections in the combination rule easy to handle. We examine a more general setting since probability mass is allocated to arbitrary subsets.
In the PLL context, we consider (Hüllermeier & Beringer, 2005; Zhang & Yu, 2015; Zhang et al., 2016; Xu et al., 2019; Wang et al., 2019) as most closely related to our proposed method as these approaches consider an instance’s neighborhood to predict its class label. To the best of our knowledge, we are the first to propose an extension of the $k$-NN classifier that keeps uncertainty from (i) and (ii) separate by leveraging Dempster-Shafer theory in the PLL context.
4 Our Method: DST-PLL
Algorithm 1 outlines our proposed method Dst-Pll, which we summarize in the following. We denote by $\operatorname{NN}_{k}(\tilde{x})\subseteq\operatorname{\mathcal{X}}\times 2^{\operatorname{\mathcal{Y}}}$ the set of the $k$-nearest neighbors of instance $\tilde{x}$ with their associated candidate labels. To predict the class label of an instance $\tilde{x}$, the algorithm first transforms information from $\tilde{x}$’s neighbors $\operatorname{NN}_{k}(\tilde{x})$ into bpas $\operatorname{m}_{i}$ (Lines 3–7), collects these into evidence set $\mathcal{E}$ (Line 8), and combines the bpas into $\operatorname{\tilde{m}}$ using Yager’s rule (Yager, 1987a, b) (Line 10). Section 4.1 elaborates on these steps and Section 4.2 details how we extract a prediction (Line 11). The latter also shows that the resulting classification rule is risk-consistent. Section 4.3 elaborates on how we extract our reject option from $\operatorname{\tilde{m}}$ (Line 12). We analyze our algorithm’s runtime in Section 4.4.
Algorithm 1 DST-PLL (Our proposed method)
0:   Partially-labeled dataset $\mathcal{D}=\{(x_{i},s_{i})\in\operatorname{\mathcal{X}}\times 2^{\operatorname{\mathcal{Y}}}\mid i\in[n]\}$, number of nearest neighbors $k$, instance $\tilde{x}\in\operatorname{\mathcal{X}}$ for inference with candidate labels $\tilde{s}\subseteq\operatorname{\mathcal{Y}}$ ($\tilde{s}=\operatorname{\mathcal{Y}}$ if $\tilde{x}$ is an unseen test instance);
0:   Prediction $g(\tilde{x})$ and reject option $\Gamma_{g}(\tilde{x})$ for instance $\tilde{x}$;
1:  $\mathcal{E}\leftarrow\emptyset$
2:  for $(x_{i},s_{i})\in\operatorname{NN}_{k}(\tilde{x})$ do
3:     if $\tilde{s}\subseteq s_{i}$ or $\tilde{s}\cap s_{i}=\emptyset$ then
4:        $\operatorname{m}_{i}:2^{\operatorname{\mathcal{Y}}}\rightarrow[0,1],\ A\mapsto\begin{cases}1&\text{if $A=\tilde{s}$,}\\ 0&\text{else}\end{cases}$
5:     else
6:        $\operatorname{m}_{i}:2^{\operatorname{\mathcal{Y}}}\rightarrow[0,1],\ A\mapsto\begin{cases}1/2&\text{if $A=\tilde{s}$ or $A=\tilde{s}\cap s_{i}$,}\\ 0&\text{else}\end{cases}$
7:     end if
8:     $\mathcal{E}\leftarrow\mathcal{E}\cup\{\operatorname{m}_{i}\}$
9:  end for
10:  $\operatorname{\tilde{m}}\leftarrow\operatorname{yager\_combination}(\tilde{s},\,\mathcal{E})$
11:  $g(\tilde{x})\leftarrow\begin{cases}\arg\max_{y\in\tilde{s}}\,\operatorname{\tilde{m}}(\{y\})&\text{if}\ \max_{y\in\tilde{s}}\,\operatorname{\tilde{m}}(\{y\})>0\text{,}\\ \text{Randomly pick from}\ \arg\max_{A\subseteq\tilde{s}}\,\operatorname{\tilde{m}}(A)&\text{else;}\end{cases}$
[…]
13:  return $(g(\tilde{x}),\,\Gamma_{g}(\tilde{x}))$
4.1 Evidence from Partially-labeled Instances
Basic probability assignments. Following the standard assumption that neighboring instances in feature space are also close in label space, we combine the evidence from the $k$ nearest neighbors $(x_{i},s_{i})\in\operatorname{NN}_{k}(\tilde{x})$ of a given instance $\tilde{x}\in\operatorname{\mathcal{X}}$ with its candidate labels $\tilde{s}\subseteq\operatorname{\mathcal{Y}}$ ($\tilde{s}=\operatorname{\mathcal{Y}}$ if $\tilde{x}$ is a test instance). For fixed $i\in[k]$, the candidate labels $s_{i}$ do not provide any valuable information if they support all ($\tilde{s}\subseteq s_{i}$) or none ($\tilde{s}\cap s_{i}=\emptyset$) of the labels in $\tilde{s}$; we use a bpa of $\operatorname{m}_{i}(\tilde{s})=1$ (Line 4). We set $\operatorname{m}_{i}(A)=1/2$ if $A=\tilde{s}$ or $A=\tilde{s}\cap s_{i}$, else $\operatorname{m}_{i}(A)=0$ (Line 6), where $1/2$ equally weights evidence. We later elaborate further on this. Note that we make the common assumption that the true label of instance $\tilde{x}$ is always in $\tilde{s}$ (Cour et al., 2011; Liu & Dietterich, 2012; Lv et al., 2020; Ni et al., 2021). While our definition of the $\operatorname{m}_{i}$-s is similar to (Denoeux, 1995), we target a more general setting. Our focal sets can be arbitrary subsets instead of only singletons or the full label set.
The bpa $\operatorname{m}_{i}$ has the following four effects on belief and plausibility as defined in (2): (i) A set of candidates $A$ has maximal belief, that is, $\operatorname{bel}_{\operatorname{m}_{i}}(A)=1$, if it covers $\tilde{s}$, that is, $\tilde{s}\subseteq A$. (ii) A set of candidates $A$ is plausible, that is, $\operatorname{pl}_{\operatorname{m}_{i}}(A)>0$, if it supports at least one of the candidate labels in $\tilde{s}$, that is, $A\cap\tilde{s}\neq\emptyset$. (iii) There is a gap, that is, $\operatorname{bel}_{\operatorname{m}_{i}}(A)<\operatorname{pl}_{\operatorname{m}_{i}}(A)$, if $A$ supports some candidate in $\tilde{s}\cap s_{i}$ but does not cover all candidates in $\tilde{s}\cap s_{i}$ or supports some candidate in $\tilde{s}$ but does not cover all candidates of $\tilde{s}$. (iv) Class labels $y\in\tilde{s}\cap s_{i}$ are maximally plausible, that is, $\operatorname{pl}_{\operatorname{m}_{i}}(\{y\})=1$.
Evidence weighting. Our definition of the $\operatorname{m}_{i}$-s (Algorithm 1, Line 6) also permits a more general view, that is, $\operatorname{m}_{i}(A)=\alpha$ if $A=\tilde{s}$, $\operatorname{m}_{i}(A)=1-\alpha$ if $A=\tilde{s}\cap s_{i}$, and $\operatorname{m}_{i}(A)=0$ otherwise, for some $\alpha\in(0,1)$. However, without further assumptions, one cannot know how relevant the information from a particular neighbor is. The setting of $\alpha=1/2$, which we use, weights supporting and conflicting evidence of all neighbors equally. In other words, if a neighbor’s evidence excludes some candidate labels from consideration, it is of equal importance compared to supporting some candidate labels. Therefore, we set $\alpha=1/2$.
Evidence combination. Given the set =$\mathcal{E}=\{\operatorname{m}_{i}\mid i\in[k]\}$, we combine all $\operatorname{m}_{i}$-s using Yager’s rule (Yager, 1987a, b). Dempster’s original rule (Dempster, 1967) enforces $\operatorname{\tilde{m}}(\emptyset)=0$ by normalization, which is criticized for its unintuitive results when facing high conflict (Zadeh, 1984). Instead, Yager’s rule first collects overlapping evidence in $\operatorname{q}:2^{\operatorname{\mathcal{Y}}}\rightarrow[0,1]$ and creates a valid bpa $\operatorname{\tilde{m}}:2^{\operatorname{\mathcal{Y}}}\rightarrow[0,1]$ by
$\operatorname{q}(A):=\sum_{\begin{subarray}{c}A_{1},\ldots,A_{k}\subseteq\operatorname{\mathcal{Y}}\\ \bigcap_{i=1}^{k}A_{i}=A\end{subarray}}\ \prod_{j=1}^{k}\operatorname{m}_{j}(A_{j})\qquad\text{and}\qquad\operatorname{\tilde{m}}(A):=\begin{cases}0&\text{if $A=\emptyset$,}\\ \operatorname{q}(\operatorname{\mathcal{Y}})+\operatorname{q}(\emptyset)&\text{if $A=\operatorname{\mathcal{Y}}$,}\\ \operatorname{q}(A)&\text{else.}\end{cases}$ (4)
We implement this efficiently with hash maps storing only the focal sets.1
4.2 Classification Rule and Consistency
After the combination into $\operatorname{\tilde{m}}$ by (4), we extract a prediction $g(\tilde{x})$ for instance $\tilde{x}$ (Line 11). We predict the class label with the highest probability mass $\arg\max_{y\in\tilde{s}}\,\operatorname{\tilde{m}}(\{y\})$ if any has non-zero mass, or else randomly pick from the subset with the most mass $\arg\max_{A\subseteq\tilde{s}}\,\operatorname{\tilde{m}}(A)$. When tied, we use the subset with the smallest cardinality.
Our classification rule yields a risk-consistent classifier, which we demonstrate in the following. As is common in the literature (Cour et al., 2011; Liu & Dietterich, 2012; Feng et al., 2020; Lv et al., 2020), we fix a label (noise) distribution (Assumption 4.1) that permits further analysis of the proposed algorithm. Appendix D experimentally verifies that Assumption 4.1 is satisfied on real-world datasets.
Assumption 4.1.
Let $\tilde{x}\in\operatorname{\mathcal{X}}$ be the instance of interest with hidden true label $\tilde{y}\in\operatorname{\mathcal{Y}}$ and $l=|\operatorname{\mathcal{Y}}|\geq 3$ classes. Its $k$ partially-labeled neighbors are $(x_{i},s_{i})\in\operatorname{NN}_{k}(\tilde{x})$. Label $\tilde{y}_{\operatorname{c}}\in\operatorname{\mathcal{Y}}\setminus\{\tilde{y}\}$ denotes the class label that co-occurs most frequently with label $\tilde{y}$ in $\tilde{x}$’s neighborhood. We assume that the true label dominates the neighborhood, that is,
$\operatorname{\mathbb{P}}(S=\underbrace{s}_{\parbox{14.22636pt}{ \tiny\centering\mbox{$=\ \{y\}\ \cup\ z$} \mbox{\text{Partitioning valid in cases $(ii)$\,--\,$(iv)$.}} \@add@centering}}\,,Y=y\mid X=x_{i})=\begin{cases}0&\text{if $s=\operatorname{\mathcal{Y}}$, $s=\emptyset$, $y\notin s$, $y\in z$, or $\tilde{y}\in z$,}\ {\hskip 7.0pt(i)}\\ \frac{1}{2^{l-2}-1}\,p_{1}&\text{if $y=\tilde{y}$, $\tilde{y}_{\operatorname{c}}\in z$, and $s\neq\operatorname{\mathcal{Y}}$,}\ {\hskip 49.57pt(ii)}\\ \frac{1}{2^{l-2}}\,p_{2}&\text{if $y=\tilde{y}$ and $\tilde{y}_{\operatorname{c}}\notin z$,}\ {\hskip 81.125pt(iii)}\\ \frac{1}{2^{l-1}-1}\frac{1}{|s|}\,p_{3}&\text{if $y\neq\tilde{y}$, $\tilde{y}\notin z$, $y\notin z$, and $s\neq\emptyset$,}\ {\hskip 27.6pt(iv)}\\ \end{cases}$
with $p_{1},p_{2},p_{3}\in(0,1)$, $p_{1}+p_{2}+p_{3}=1$, and $p_{1}\geq p_{2}\geq p_{3}>0$.
Because $p_{1}+p_{2}>p_{3}$, Assumption 4.1 implies $\operatorname{\mathbb{P}}_{XY}(Y=\tilde{y}\mid X=x_{i})>\operatorname{\mathbb{P}}_{XY}(Y\neq\tilde{y}\mid X=x_{i})$, that is, points that are close in feature space are likely to have similar class labels. In the following, we elaborate on the four cases $(i)$ – $(iv)$.
  • If a candidate set contains all labels ($s=\operatorname{\mathcal{Y}}$), no labels ($s=\emptyset$), or does not allow to be partitioned into true label $y$ and false-positive labels $z$, it is not a valid candidate set (Hüllermeier & Beringer, 2005; Cour et al., 2011; Liu & Dietterich, 2012; Zhang & Yu, 2015). $(i)$ handles this pathological setting by assigning zero probability.
  • In $(ii)$, $\tilde{x}$’s label $\tilde{y}$ is also the label of neighbor $x_{i}$ and $\tilde{y}_{\operatorname{c}}$, with which $\tilde{y}$ is most commonly confused, is also part of the candidate set $s_{i}$. There are $12^{l-2}-$ 1 sets $s$ that contain $\tilde{y}$ and $\tilde{y}_{\operatorname{c}}$ and are not $\operatorname{\mathcal{Y}}$.
  • In $(iii)$, $\tilde{x}$’s label $\tilde{y}$ is also the label of neighbor $x_{i}$. Label $\tilde{y}_{\operatorname{c}}$ is not part of the candidate set $s_{i}$ of neighbor $x_{i}$. There are $2^{l-2}$ sets $s$ that contain $\tilde{y}$ but not $\tilde{y}_{\operatorname{c}}$.
  • In $(iv)$, neighbor $x_{i}$ has different labels altogether and is irrelevant for explaining the class label of instance $\tilde{x}$. There are $12^{l-1}-$ 1 sets that do not contain the label $\tilde{y}$ (excluding the empty set). Each of the $|s|$ candidates can be the neighbor’s correct label.
Theorem 4.3, which establishes risk-consistency, hinges on the following intermediate results in Lemma 4.2. The main difficulty arises in the application of Yager’s rule (4) to produce $\operatorname{\tilde{m}}$ as the focal sets of the $\operatorname{m}_{i}$-s can be arbitrary subsets. The proof of Lemma 4.2 provides a closed-form solution.
Lemma 4.2.
Assume the setting of Assumption 4.1, let $\tilde{x}\in\operatorname{\mathcal{X}}$ be the instance of interest, $\tilde{y}\in\operatorname{\mathcal{Y}}$ its hidden true label, $\tilde{y}_{\operatorname{c}}\in\operatorname{\mathcal{Y}}\setminus\{\tilde{y}\}$ the incorrect label with which $\tilde{y}$ is most often confused, and $\operatorname{\tilde{m}}$ the resulting probability mass as determined by Algorithm 1. Then, the following hold:
  1. (i)
    $\mathbb{E}_{\operatorname{\mathbb{P}}}[\operatorname{\tilde{m}}(\{\tilde{y}\})\mid X=\tilde{x}]>0$, and
  2. (ii)
    $\mathbb{E}_{\operatorname{\mathbb{P}}}[\operatorname{\tilde{m}}(\{\tilde{y}\})\mid X=\tilde{x}]>\mathbb{E}_{\operatorname{\mathbb{P}}}[\operatorname{\tilde{m}}(\{\tilde{y}_{c}\})\mid X=\tilde{x}]$.
As the probability mass of the hidden true label $\tilde{y}$ is positive in expectation by $(i)$, we can reduce our analysis to the first case of our decision rule (Algorithm 1, Line 11). (ii) shows how Assumption 4.1 propagates when applying Yager’s rule on all $k$ neighbors’ $\operatorname{m}_{i}$-s. The label that dominates the neighborhood obtains the highest probability mass. To our knowledge, it is the first time that such a result has been shown for arbitrary focal sets in $\operatorname{\tilde{m}}$. Denoeux (1995) analyze the special case when all focal sets are singletons or the full label space.
To put Theorem 4.3 into context, we recall the concept of risk consistency (Devroye et al., 1996). As Assumption 4.1 fixes a distribution $\operatorname{\mathbb{P}}$, one can quantify the classification risk $\operatorname{\mathcal{R}}(g)$ of a classifier $g$ as defined in Section 2. The Bayes classifier is defined by $g^{*}=\arg\min_{g:\operatorname{\mathcal{X}}\rightarrow\operatorname{\mathcal{Y}}}\operatorname{\mathcal{R}}(g)$; it has the least overall risk. Let $g_{n}$ be the classifier trained by Alg. 1 with $n$ instances. Its empirical risk ^$\hat{\operatorname{\mathcal{R}}}(g_{n})$ can be computed by substituting the expectation with a sample mean. It is risk-consistent if ^$\hat{\operatorname{\mathcal{R}}}(g_{n})\to\operatorname{\mathcal{R}}(g^{*})$ for $n\to\infty$ almost surely. Theorem 4.3 establishes the risk consistency of the proposed classifier.
Theorem 4.3.
Assume the setting of Assumption 4.1, let $g_{n}:\operatorname{\mathcal{X}}\to\operatorname{\mathcal{Y}}$ be the classifier trained by Algorithm 1 with $n$ training instances, $g^{*}:\operatorname{\mathcal{X}}\to\operatorname{\mathcal{Y}}$ the Bayes classifier, and $\tilde{x}\in\operatorname{\mathcal{X}}$ a fixed instance with unknown true label $\tilde{y}$. Then, the following hold:
  1. (i)
[…]
  2. (ii)
    $\lim_{n\to\infty}\bigl{(}\hat{\operatorname{\mathcal{R}}}(g_{n})-\operatorname{\mathcal{R}}(g^{*})\bigr{)}=0$ almost surely.
4.3 Reject Option
To limit the impact of misclassification, our method provides a reject option $\Gamma_{g}$, that is, the algorithm can abstain from individual predictions if unsure (Algorithm 1, Line 12). Recall from (3) that belief describes the lower-bound probability and plausibility the upper-bound probability of the credal set $\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$. Let $\operatorname{\tilde{m}}$ be the resulting bpa as determined by Algorithm 1 and
$\Delta_{\operatorname{\tilde{m}}}:=\operatorname{bel}_{\operatorname{\tilde{m}}}(\{\hat{y}\})-\max_{y\in\tilde{s}\setminus\{\hat{y}\}}\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y\})\quad\text{with}\quad\hat{y}:=\arg\max_{y\in\tilde{s}}\,\operatorname{\tilde{m}}(\left\{y\right\})\text{.}$ (5)
$\Delta_{\operatorname{\tilde{m}}}$ satisfies several desirable properties: Lemma 4.4 $(i)$ states that $\Delta_{\operatorname{\tilde{m}}}\leq 0$ if the classification decision has been taken randomly (Algorithm 1, Line 11, second case). Lemma 4.4 $(ii)$ states that, if $\Delta_{\operatorname{\tilde{m}}}>0$, $g(\tilde{x})$ will always be our prediction independent of $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$. Finally, Lemma 4.4 $(iii)$ states that a mass of more than 1/2 is sufficient, but not necessary, for $\Delta_{\operatorname{\tilde{m}}}>0$.
Lemma 4.4.
Let $\operatorname{\mathcal{Y}}$ be the label space, $\tilde{x}\in\operatorname{\mathcal{X}}$ the instance of interest, $\tilde{s}\subseteq\operatorname{\mathcal{Y}}$ its candidate labels ($\tilde{s}=\operatorname{\mathcal{Y}}$ if $\tilde{x}$ is a test instance), $g(\tilde{x})$ our algorithm’s prediction, and $\operatorname{\tilde{m}}$ the resulting probability mass as determined by Algorithm 1. Then, the following hold:
  1. (i)
    If $g(\tilde{x})$ has been picked randomly (Algorithm 1, Line 11, second case), then $\Delta_{\operatorname{\tilde{m}}}\leq 0$,
  2. (ii)
    if $\Delta_{\operatorname{\tilde{m}}}>0$, then ($\operatorname{\mathbb{P}}(\{g(\tilde{x})\})>\operatorname{\mathbb{P}}(\{y\})$ for all $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ and $y\in\tilde{s}\setminus\{g(\tilde{x})\}$, and
  3. (iii)
    if $\operatorname{\tilde{m}}(\left\{g(\tilde{x})\right\})>1/2$, then $\Delta_{\operatorname{\tilde{m}}}>0$. The converse does not hold.
Based on these considerations, we define the accept and reject regions of our method accordingly.
  1. 1.
    Accept. When $\Delta_{\operatorname{\tilde{m}}}>0$, we are in the accept region, that is, $\Gamma_{g}(\tilde{x})=\{\hat{y}\}$: The lower-bound probability of the class label $\hat{y}$ is greater than the upper-bound probability of any other class label $y\neq\hat{y}$, that is, ($\operatorname{\mathbb{P}}(\{\hat{y}\})>\operatorname{\mathbb{P}}(\{y\})$ for all $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ and $y\neq\hat{y}$.
  2. 2.
    Reject. When $\Delta_{\operatorname{\tilde{m}}}\leq 0$, we are in the reject region, that is, $\Gamma_{g}(\tilde{x})=\emptyset$: The lower-bound probability of the class label $\hat{y}$ is less than or equal to the upper-bound probability of another class label $y\neq\hat{y}$. There exists $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ and $y\neq\hat{y}$ with ($\operatorname{\mathbb{P}}(\{\hat{y}\})\leq\operatorname{\mathbb{P}}(\{y\})$.
Moreover, Lemma 4.4 $(iii)$ also shows that we reject fewer predictions than the $k$-nearest neighbor reject option by Hellman (1970), which needs always more than 1/2 of all votes to be certain. Our method can accept predictions with less than 1/2 of all probability mass on a single class label, while still satisfying Lemma 4.4 $(ii)$, that is, a prediction remains unchanged independent of $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ if it is accepted.
4.4 Runtime Complexity
Algorithm 1 yields a worst-case runtime complexity of $\mathcal{O}\!\left(dkn+l\left(\min\!\left(2^{k},2^{l}\right)+\max\!\left(k,l\right)\right)\right)$. Since $k$ and $l$ are constant, the nearest-neighbor search dominates. The average total runtime of the search is $\mathcal{O}\!\left(dk\log n\right)$. An in-depth discussion can be found in Appendix A.4.
5 Experiments
Section 5.1 lists the tested methods, Section 5.2 shows our setup, and Section 5.3 collects our main findings. Additional results and a description of all hyperparameters can be found in Appendix D.
5.1 Algorithms for Comparison
While many PLL algorithms exist (see Section 3), we focus on methods commonly used in the literature. In total, we consider ten methods: Pl-Knn (Hüllermeier & Beringer, 2005), Pl-Svm (Nguyen & Caruana, 2008), Ipal (Zhang & Yu, 2015), Pl-Ecoc (Zhang et al., 2017), Proden (Lv et al., 2020), Rc (Feng et al., 2020), Cc (Feng et al., 2020), Valen (Xu et al., 2021), Pop (Xu et al., 2023), and Dst-Pll (our proposed method). We choose their parameters as recommended and use the same base models for all of the neural network approaches. As base models, we pick the LeNet architecture (LeCun et al., 1998) for the MNIST datasets and a $d$-300-300-300-$l$ MLP (Werbos, 1974) for all other datasets. For Pl-Knn and our method, we use variational auto-encoders (Kingma & Welling, 2014) to reduce the feature space dimensionality of the MNIST datasets and compute the nearest neighbors on the hidden representations. As our competitors do not provide a reject option, we use a threshold on their model’s confidences, that is, the maximum probability outputs, to evaluate the trade-off between the fraction and accuracy of non-rejected predictions. Appendix D.2 discusses the choices of all hyperparameters in more detail.
Table 1: Average test-set accuracies ($\pm$ std.) on the UCI, MNIST, and real-world datasets. The UCI and MNIST datasets are artificially augmented with class-dependent and instance-dependent noise. The best algorithms (highest accuracy) as well as algorithms with non-significant differences are emphasized. The significance analysis uses a paired student t-test with level $\alpha=0.05$.
Algorithms UCI datasets MNIST-like datasets Real-world
Class-dep. Inst.-dep. Class-dep. Inst.-dep.
Pl-Knn (2005) 81.2 ($\pm$ 13.9) 75.7 ($\pm$ 11.4) 92.2 ($\pm$ 04.1) 84.9 ($\pm$ 07.2) 53.4 ($\pm$ 11.0)
Pl-Svm (2008) 62.3 ($\pm$ 16.3) 42.5 ($\pm$ 17.1) 67.5 ($\pm$ 09.8) 46.8 ($\pm$ 06.9) 39.1 ($\pm$ 09.6)
Ipal (2015) 79.4 ($\pm$ 16.9) 75.6 ($\pm$ 18.1) 92.9 ($\pm$ 04.3) 88.5 ($\pm$ 06.6) 58.7 ($\pm$ 09.8)
Pl-Ecoc (2017) 63.7 ($\pm$ 13.1) 66.5 ($\pm$ 12.1) 64.3 ($\pm$ 14.3) 51.6 ($\pm$ 11.2) 46.2 ($\pm$ 10.2)
Proden (2020) 81.4 ($\pm$ 14.0) 78.1 ($\pm$ 13.2) 93.9 ($\pm$ 04.5) 88.3 ($\pm$ 06.1) 64.0 ($\pm$ 07.9)
Rc (2020) 82.0 ($\pm$ 13.9) 78.3 ($\pm$ 13.6) 94.0 ($\pm$ 04.1) 88.0 ($\pm$ 06.2) 61.5 ($\pm$ 10.0)
Cc (2020) 81.8 ($\pm$ 14.2) 78.5 ($\pm$ 13.8) 94.0 ($\pm$ 04.5) 89.6 ($\pm$ 05.7) 49.0 ($\pm$ 29.6)
Valen (2021) 79.6 ($\pm$ 15.3) 75.3 ($\pm$ 12.8) 91.9 ($\pm$ 04.1) 82.9 ($\pm$ 08.9) 63.3 ($\pm$ 09.5)
Pop (2023) 81.7 ($\pm$ 13.9) 78.0 ($\pm$ 13.1) 93.8 ($\pm$ 04.5) 88.0 ($\pm$ 06.1) 64.0 ($\pm$ 08.2)
Dst-Pll (ours) 80.7 ($\pm$ 14.1) 75.2 ($\pm$ 10.9) 92.0 ($\pm$ 04.2) 84.6 ($\pm$ 07.1) 52.1 ($\pm$ 11.6)
Table 2: Average empirical reject option risk ^$\hat{\operatorname{\mathcal{R}}}_{\lambda}(\Gamma_{g})$ as defined in (1) for various values of $\lambda$. The best algorithms (smallest risk) as well as algorithms with non-significant differences are emphasized. The significance analysis uses a paired student t-test with level $\alpha=0.05$.
Algorithms Average ^$\hat{\operatorname{\mathcal{R}}}_{\lambda}(\Gamma_{g})$ ($\pm$ std.) across all experimental settings
$\lambda=0.00$ $\lambda=0.05$ $\lambda=0.10$ $\lambda=0.15$ $\lambda=0.20$
Pl-Knn (2005) 0.10 ($\pm$ 0.14) 0.13 ($\pm$ 0.16) 0.17 ($\pm$ 0.17) 0.20 ($\pm$ 0.18) 0.24 ($\pm$ 0.19)
Pl-Svm (2008) 0.21 ($\pm$ 0.29) 0.25 ($\pm$ 0.30) 0.29 ($\pm$ 0.31) 0.32 ($\pm$ 0.32) 0.36 ($\pm$ 0.33)
Ipal (2015) 0.19 ($\pm$ 0.17) 0.20 ($\pm$ 0.18) 0.21 ($\pm$ 0.19) 0.22 ($\pm$ 0.20) 0.23 ($\pm$ 0.21)
Pl-Ecoc (2017) 0.26 ($\pm$ 0.28) 0.31 ($\pm$ 0.28) 0.35 ($\pm$ 0.28) 0.40 ($\pm$ 0.28) 0.45 ($\pm$ 0.28)
Proden (2020) 0.12 ($\pm$ 0.11) 0.13 ($\pm$ 0.11) 0.14 ($\pm$ 0.12) 0.15 ($\pm$ 0.12) 0.16 ($\pm$ 0.13)
Rc (2020) 0.14 ($\pm$ 0.14) 0.14 ($\pm$ 0.14) 0.15 ($\pm$ 0.14) 0.16 ($\pm$ 0.14) 0.17 ($\pm$ 0.14)
Cc (2020) 0.16 ($\pm$ 0.18) 0.16 ($\pm$ 0.19) 0.17 ($\pm$ 0.20) 0.18 ($\pm$ 0.20) 0.19 ($\pm$ 0.21)
Valen (2021) 0.20 ($\pm$ 0.14) 0.20 ($\pm$ 0.14) 0.20 ($\pm$ 0.14) 0.21 ($\pm$ 0.14) 0.21 ($\pm$ 0.14)
Pop (2023) 0.12 ($\pm$ 0.11) 0.13 ($\pm$ 0.11) 0.14 ($\pm$ 0.12) 0.15 ($\pm$ 0.12) 0.16 ($\pm$ 0.13)
Dst-Pll (ours) 0.05 ($\pm$ 0.07) 0.07 ($\pm$ 0.08) 0.10 ($\pm$ 0.10) 0.12 ($\pm$ 0.11) 0.15 ($\pm$ 0.12)
(a) Ecoli (inst.-dep. noise)    (b) KMNIST (inst.-dep. noise)        (c) msrc-v2
[…]
5.2 Experimental Setup
Data. Following the default protocol (Cour et al., 2011; Zhang & Yu, 2015; Lv et al., 2020; Xu et al., 2023), we conduct several experiments using datasets for supervised learning with added artificial noise as well as experiments on real-world partially-labeled data. We repeat all experiments five times to report averages and standard deviations. For the supervised datasets, we use the ecoli (Horton & Nakai, 1996), multiple-features (Duin, 2002), pen-digits (Alpaydin & Alimoglu, 1998), semeion (Buscema & Terzi, 2008), solar-flare (Dodson & Hedeman, 1989), statlog-landsat (Srinivasan, 1993), and theorem datasets (Bridge et al., 2013) from the UCI repository (Bache & Lichman, 2013). Also, we use the MNIST (LeCun et al., 1999), KMNIST (Clanuwat et al., 2018), and FMNIST datasets (Xiao et al., 2017). For the partially labeled data, we use the bird-song (Briggs et al., 2012), flickr (Huiskes & Lew, 2008), yahoo-news (Guillaumin et al., 2010), and msrc-v2 datasets (Liu & Dietterich, 2012).
[…]
5.3 Results
Prediction Performance. Table 1 shows the average test-set accuracies and standard deviations over all UCI datasets with class- and instance-dependent noise, MNIST datasets with class- and instance-dependent noise, and the real-world datasets. The algorithms with the highest accuracies as well as the algorithms with non-significant differences using a paired t-test with level $\alpha=0.05$ are emphasized. Our approach (Dst-Pll) performs comparably to the other methods. We note that none of the methods is best across all settings. For example, Cc performs best in four out of five settings but is significantly outperformed by our approach on the real-world experiments.
[…]
Table 2 shows the average empirical reject-option risk and standard deviation across all experiments for varying $\lambda$. We compute ^$\hat{\operatorname{\mathcal{R}}}_{\lambda}(\Gamma_{g})=\hat{\operatorname{err}}(\Gamma_{g})+\lambda\hat{\operatorname{r}}(\Gamma_{g})$ by using ground-truth information to calculate the non-reject error $\hat{\operatorname{err}}(\Gamma_{g})$ and counting the number of rejects to calculate the reject rate $\hat{\operatorname{r}}(\Gamma_{g})$. The algorithms with the lowest risks as well as the algorithms with non-significant differences using a paired t-test with level $\alpha=0.05$ are emphasized. When misclassification is costly ($\lambda\leq 0.2$), our method provides the significantly best trade-off compared to our competitors. In contrast, when rejecting predictions is costly ($\lambda>0.2$), the methods in Table 1 are to be preferred.
Figure 1 shows the reject trade-off for varying confidence (0 to 1) and $\Delta_{\operatorname{\tilde{m}}}$ (-1 to 1) thresholds on the ecoli and KMNIST datasets with instance-dependent noise as well as on the msrc-v2 real-world dataset. The x-axes show the fractions of predictions that are rejected. The y-axes show the accuracies of predictions that are not rejected. Most methods have monotonical growth: The more predictions are rejected, the more accurate are non-rejected predictions. Also, it is desirable to be close to the top-left corner of the plots as one wants to achieve high accuracy by rejecting as few predictions as possible. Our method provides the significantly best trade-offs compared to competitors (Table 2). These results empirically demonstrate our method’s superiority in reliably rejecting unsure predictions.
6 Conclusions
We addressed the PLL problem. Current methods offer good predictive performance but even the best algorithms give incorrect predictions. Therefore, we discussed the option to abstain from predictions, that is, to reject individual predictions if unsure. The proposed approach offers the best trade-off between the number and accuracy of non-rejected predictions compared to competitors. Experiments demonstrate the effectiveness of our algorithm. Further, we prove our method’s risk consistency.
Acknowledgments and Disclosure of Funding
Acknowledgements. TBD.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our claims accurately reflect the paper’s contributions. Section 2 defines the problem, Section 4 presents our method and its analysis, and Section 5 shows our empirical findings.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We explicitly state our assumptions in Section 4.1 and Section 4.2. Our consistency analysis is limited to the concrete setting considered in Assumption 4.1. We discuss the computational efficiency of our proposed algorithm in Section 4.4.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: To the best of our knowledge, we state all relevant prerequisites of the theorems and lemmas as well as give full proofs in Appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 5.2 describes our experimental setup. We make all our code and data publicly available with instructions on how to run our experiments. [^3] Repository: https://github.com/anon1248/pll-with-a-reject-option. Full source code and data are also available in the supplemental material.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We make all our code and data publicly available under the MIT license.3 The readme contains all relevant instructions to reproduce the experiments including instructions for Linux, MacOS, and Windows operating systems. We include scripts to regenerate all plots and tables of the paper.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Section 5.1 and Appendix D.2 provide a detailed list of all hyperparameters and base models used in our experiments. Further details are provided in our linked repository’s readme.3
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Table 1 and Table 2 include, both, average values and standard deviations as well as highlight significant differences using a paired student t-test with level $\alpha=0.05$. Figure 1 shows examples of the performance of our reject option that are also included in the values of Table 2.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Appendix D.2 describes the compute resources needed as well as our machine used. Further details are provided in our linked repository’s readme.3
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We reviewed the NeurIPS Code of Ethics and conform to all its elements.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: As this paper presents foundational research, we believe that it does not have a direct societal impact or the impact is quite limited.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We believe that the paper does not pose such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We properly cite all used datasets in Section 5.2. Furthermore, we include all relevant third-party licenses in our published repository.3
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We make all of our code and data publicly available.3 We describe our algorithm in Section 4, the used data in Section 5.2 as well as provide a readme with the repository containing both.3
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our research does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our research does not involve crowdsourcing or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
Appendix A Proofs
This section collects our proofs. The proof of Lemma 4.2 is in Appendix A.1. We prove our main statement, Theorem 4.3, in Appendix A.2. Appendix A.3 contains the proof of Lemma 4.4.
A.1 Proof of Lemma 4.2
Part $(i)$. The expected probability mass of $\tilde{y}$ is given by (27). The non-negativity of the binomial coefficients implies that it is sufficient to show that
$\binom{k-h}{i}\prod_{a=1}^{l-2}\binom{k-h}{j_{a}}\geq\sum_{b=1}^{\min\left(i,j_{1},\dots,j_{l-2}\right)}\left(-1\right)^{b+1}\binom{k-h}{b}\binom{k-h-b}{i-b}\prod_{c=1}^{l-2}\binom{k-h-b}{j_{c}-b}\text{,}$ (6)
where $k\in\mathbb{N}$, $h,i,j_{a}\in\mathbb{N}_{0}$ with $0\leq i<k-h$ and $0\leq j_{a}<k-h$ with $a\in[l-2]$, and that the inequality is strict in at least one case.
The proof of (6) proceeds by induction over $k$. Note that the choices of $i$ and $j_{a}$ are interchangeable (see Appendix B). Hence, w.l.o.g., we assume for the remainder of this proof that $i\leq j_{a}$ for all $a\in\left[l-2\right]$. Further, we set $k^{\prime}=k-h\in\mathbb{N}$.
Proposition. For $k^{\prime}\in\mathbb{N}$, it holds that
$\binom{k^{\prime}}{i}\prod_{a=1}^{l-2}\binom{k^{\prime}}{j_{a}}\geq\sum_{b=1}^{i}\left(-1\right)^{b+1}\binom{k^{\prime}}{b}\binom{k^{\prime}-b}{i-b}\prod_{c=1}^{l-2}\binom{k^{\prime}-b}{j_{c}-b}\text{,}$ (7)
with $0\leq i\leq j_{a}<k^{\prime}$ for all $a\in\left[l-2\right]$.
Base case. For $k^{\prime}=1$, it holds that $h=i=j_{a}=0$ ($a\in\left[l-2\right]$). Then, all binomial coefficients on the l.h.s. are $1$. The sum on the r.h.s. is empty and thus 00. Hence, l.h.s. $>$ r.h.s. ($1>0$), which also renders (27) strictly positive.
Induction step. Let $k^{\prime}=k^{\prime}+1$. The l.h.s. yields
$\ \binom{k^{\prime}+1}{i}\prod_{a=1}^{l-2}\binom{k^{\prime}+1}{j_{a}}\stackrel{{\scriptstyle(i)}}{{=}}\left(\frac{k^{\prime}+1}{k^{\prime}+1-i}\prod_{a=1}^{l-2}\frac{k^{\prime}+1}{k^{\prime}+1-j_{a}}\right)\binom{k^{\prime}}{i}\prod_{a=1}^{l-2}\binom{k^{\prime}}{j_{a}}$ (8)
$\stackrel{{\scriptstyle(ii)}}{{\geq}}$ $\ \left(\frac{k^{\prime}+1}{k^{\prime}+1-i}\prod_{a=1}^{l-2}\frac{k^{\prime}+1}{k^{\prime}+1-j_{a}}\right)\sum_{b=1}^{i}\left(-1\right)^{b+1}\binom{k^{\prime}}{b}\binom{k^{\prime}-b}{i-b}\prod_{c=1}^{l-2}\binom{k^{\prime}-b}{j_{c}-b}$ (9)
=$\stackrel{{\scriptstyle(iii)}}{{=}}$ $\ \sum_{b=1}^{i}\left(-1\right)^{b+1}\frac{k^{\prime}+1}{k^{\prime}+1-i}\binom{k^{\prime}}{b}\binom{k^{\prime}-b}{i-b}\prod_{c=1}^{l-2}\frac{k^{\prime}+1}{k^{\prime}+1-j_{c}}\binom{k^{\prime}-b}{j_{c}-b}$ (10)
=$\stackrel{{\scriptstyle(iv)}}{{=}}$ $\ \sum_{b=1}^{i}\left(-1\right)^{b+1}\underbrace{\frac{k^{\prime}+1}{k^{\prime}+1-b}\binom{k^{\prime}}{b}}_{\stackrel{{\scriptstyle(i^{\prime})}}{{=}}\binom{k^{\prime}+1}{b}}\underbrace{\frac{k^{\prime}+1-b}{k^{\prime}+1-i}\binom{k^{\prime}-b}{i-b}}_{\stackrel{{\scriptstyle(i^{\prime})}}{{=}}\binom{k^{\prime}+1-b}{i-b}}\prod_{c=1}^{l-2}\frac{k^{\prime}+1}{k^{\prime}+1-b}\underbrace{\frac{k^{\prime}+1-b}{k^{\prime}+1-j_{c}}\binom{k^{\prime}-b}{j_{c}-b}}_{\stackrel{{\scriptstyle(i^{\prime})}}{{=}}\binom{k^{\prime}+1-b}{j_{c}-b}}$ (11)
=$\stackrel{{\scriptstyle(v)}}{{=}}$ $\ \sum_{b=1}^{i}\left(-1\right)^{b+1}\left(\frac{k^{\prime}+1}{k^{\prime}+1-b}\right)^{l-2}\binom{k^{\prime}+1}{b}\binom{k^{\prime}+1-b}{i-b}\prod_{c=1}^{l-2}\binom{k^{\prime}+1-b}{j_{c}-b}$ (12)
$\stackrel{{\scriptstyle(vi)}}{{\geq}}$ $\ \sum_{b=1}^{i}\left(-1\right)^{b+1}\binom{k^{\prime}+1}{b}\binom{k^{\prime}+1-b}{i-b}\prod_{c=1}^{l-2}\binom{k^{\prime}+1-b}{j_{c}-b}\text{,}$ (13)
where $(i)$ holds as
$\binom{k^{\prime}+1}{i}=\frac{(k^{\prime}+1)!}{i!(k^{\prime}+1-i)!}=\frac{k^{\prime}+1}{k^{\prime}+1-i}\frac{k^{\prime}!}{i!(k^{\prime}-i)!}=\frac{k^{\prime}+1}{k^{\prime}+1-i}\binom{k^{\prime}}{i},$
the proposition yields $(ii)$, in $(iii)$ we rearrange the sum and collect both products, and $(iv)$ expands the fractions. In $(v)$, we rearrange the factor within the product and apply the $(i^{\prime})$-s, which hold by similar calculations as $(i)$, and Lemma B.1 with $k=k^{\prime}+1$ implies $(vi)$. These derivations and the observation that the base case guarantees the inequality to be strict conclude the proof.
Part $(ii)$. By comparing the closed-form expressions (with $m=i=t$) of $\tilde{y}$ and $\tilde{y}_{\operatorname{c}}$ given in (27) and  (32), respectively, we obtain
$\ \mathbb{E}_{\operatorname{\mathbb{P}}}\!\left[\operatorname{\tilde{m}}(\left\{\tilde{y}\right\})\mid X=\tilde{x}\right]>\mathbb{E}_{\operatorname{\mathbb{P}}}\!\left[\operatorname{\tilde{m}}(\{\tilde{y}_{\operatorname{c}}\})\mid X=\tilde{x}\right]$ (14)
$\Leftrightarrow$ $\ \left(\frac{1}{2^{l-2}}p_{2}\right)^{k-h-m}>\left(\frac{1}{2^{l-1}-1}p_{3}\right)^{k-h-m}$ (15)
$\Leftrightarrow$ $\ \frac{1}{2^{l-2}}p_{2}>\frac{1}{2^{l-1}-1}p_{3}\ \stackrel{{\scriptstyle(i)}}{{\Leftrightarrow}}\ \frac{2^{l-1}-1}{2^{l-2}}>1\ \Leftrightarrow\ 2^{l-2}>1\text{,}$ (16)
where $(i)$ holds as Assumption 4.1 guarantees that $p_{2}\geq p_{3}>0$. The last statement is satisfied for $l\geq 3$, which shows that the dominance assumption w.r.t. the label distribution propagates through the computation of belief when using Yager’s rule. We note that the result is independent of $p_{1}$ as a candidate set can not distinguish the belief in $\tilde{y}$ and $\tilde{y}_{\operatorname{c}}$ when these co-occur.
A.2 Proof of Theorem 4.3
With Assumption 4.1, Theorem 4.3 is a consequence of Lemma 4.2. We detail this in the following.
[…]
$\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}}[\hat{\operatorname{\mathcal{R}}}(g_{n})]=\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}}\!\left[\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathcal{L}}(g_{n}(x_{i}),s_{i})\right]=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}}[\operatorname{\mathcal{L}}(g_{n}(x_{i}),s_{i})]\overset{(i)}{=}0\text{,}$ (17)
which is zero by part $(i)$ of the statement as, in expectation, we consistently predict the true label $y_{i}$ for any fixed label $x_{i}$.
By the law of large numbers, we have
$\lim_{n\to\infty}\hat{\operatorname{\mathcal{R}}}(g_{n})=\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathcal{L}}(g_{n}(x_{i}),s_{i})=0\ \ \text{almost surely,}$, (18)
because $\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}}[\hat{\operatorname{\mathcal{R}}}(g)]=0$ as stated above. As $\operatorname{\mathcal{R}}\!\left(g^{*}\right)\geq 0$, we conclude
$0\leq\lim_{n\to\infty}\bigl{(}\hat{\operatorname{\mathcal{R}}}(g_{n})-\operatorname{\mathcal{R}}(g^{*})\bigr{)}\leq\lim_{n\to\infty}\hat{\operatorname{\mathcal{R}}}(g_{n})=0\ \ \text{almost surely,}$, (19)
which establishes $(ii)$.
A.3 Proof of Lemma 4.4
[…]
$\Delta_{\operatorname{\tilde{m}}}=\operatorname{bel}_{\operatorname{\tilde{m}}}(\{y\})-\max_{y^{\prime}\in\tilde{s}\setminus\{y\}}\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y^{\prime}\})=0-\max_{y^{\prime}\in\tilde{s}\setminus\{y\}}\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y^{\prime}\})\leq 0\text{,}$ (20)
independent of the choice of $y\in\operatorname{\mathcal{Y}}$. Therefore, $\Delta_{\operatorname{\tilde{m}}}\leq 0$.
Part $(ii)$. Given an instance $\tilde{x}\in\operatorname{\mathcal{X}}$ with its candidate set $\tilde{s}\subseteq\operatorname{\mathcal{Y}}$ and its associated prediction $g(\tilde{x})$ as described in Algorithm 1, we assume that $\Delta_{\operatorname{\tilde{m}}}>0$. By the contraposition of part $(i)$, $g(\tilde{x})=\hat{y}$ with $\hat{y}=\arg\max_{y\in\tilde{s}}\,\operatorname{\tilde{m}}(\left\{y\right\})$. From (5) and $\Delta_{\operatorname{\tilde{m}}}>0$, it follows that $\operatorname{bel}_{\operatorname{\tilde{m}}}(\{\hat{y}\})>\max_{y\in\tilde{s}\setminus\{\hat{y}\}}\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y\})$, which is equivalent to $\operatorname{bel}_{\operatorname{\tilde{m}}}(\{\hat{y}\})>\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y\})$ for all $y\in\tilde{s}\setminus\{\hat{y}\}$. For all $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$, it holds that $\operatorname{bel}_{\operatorname{\tilde{m}}}\!(A)\leq\operatorname{\mathbb{P}}(A)\leq\operatorname{pl}_{\operatorname{\tilde{m}}}(A)$ for all $A\subseteq\operatorname{\mathcal{Y}}$ by (3). Therefore, ($\operatorname{\mathbb{P}}(\{y\})\leq\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y\})<\operatorname{bel}_{\operatorname{\tilde{m}}}(\{\hat{y}\})\leq\operatorname{\mathbb{P}}(\{\hat{y}\})$ for all $y\in\tilde{s}\setminus\{\hat{y}\}$.
Part $(iii)$. ($\Rightarrow$): Given an instance $\tilde{x}\in\operatorname{\mathcal{X}}$ with its candidate set $\tilde{s}\subseteq\operatorname{\mathcal{Y}}$ and its associated prediction $g(\tilde{x})$ as described in Algorithm 1, we assume that $\operatorname{\tilde{m}}(\{g(\tilde{x})\})>1/2$. Because $\max_{y\in\tilde{s}}\,\operatorname{\tilde{m}}(\left\{y\right\})>0$, we are in the first case in Line 11 of Algorithm 1. Therefore, $g(\tilde{x})=\hat{y}$ with $\hat{y}=\arg\max_{y\in\tilde{s}}\,\operatorname{\tilde{m}}(\left\{y\right\})$. As $\sum_{A\subseteq\operatorname{\mathcal{Y}}}\operatorname{\tilde{m}}\!\left(A\right)=1$ and $\operatorname{\tilde{m}}(\{\hat{y}\})>1/2$, it holds that $\sum_{A\subseteq\operatorname{\mathcal{Y}},\,A\neq\left\{\hat{y}\right\}}\operatorname{\tilde{m}}\!\left(A\right)<1/2$. Then, for all $y\in\tilde{s}$ with $y\neq\hat{y}$, $\operatorname{pl}_{\operatorname{\tilde{m}}}\!\left(\left\{y\right\}\right)=\sum_{A\subseteq\operatorname{\mathcal{Y}},\,A\cap\left\{y\right\}\neq\emptyset}\operatorname{\tilde{m}}\!\left(A\right)<1/2$. Hence, $\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y\})<\operatorname{bel}_{\operatorname{\tilde{m}}}(\{\hat{y}\})$ for all $y\in\tilde{s}$ with $y\neq\hat{y}$. Therefore, $\Delta_{\operatorname{\tilde{m}}}=\operatorname{bel}_{\operatorname{\tilde{m}}}(\{\hat{y}\})-\max_{y\in\tilde{s}\setminus\{\hat{y}\}}\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y\})>0$.
[…]
A.4 Runtime Analysis of Algorithm 1
Looking at Algorithm 1, we decompose the overall runtime of our approach into $k$-times querying one nearest neighbor and creating its bpa (Lines 2–9), the cost of Yager’s rule (Line 10), as well as extracting predictions in Lines 11-12.
Using the ball-tree data structure (Omohundro, 1989), querying one neighbor takes $\mathcal{O}\!\left(d\log n\right)$ time on average. In the worst case, query time is $\mathcal{O}\!\left(dn\right)$. One builds a ball-tree in $\mathcal{O}\!\left(dn\log n\right)$ time.
We construct a bpa $\operatorname{m}_{i}$ by storing its focal sets within a hash map and combine all $\operatorname{m}_{i}$-s as defined in (4). There are at most $\min\!\left(2^{k},2^{l}\right)$ focal sets of $\operatorname{\tilde{m}}$: Each $\operatorname{m}_{i}$ has at most two focal sets producing $2^{k}$ combinations and there are at most $2^{l}$ possible subsets of $\operatorname{\mathcal{Y}}$. We take the minimum as both are upper bounds. Looking up a focal set in the hash-map requires $\mathcal{O}\!\left(l\right)$ time. Extracting a prediction and the reject option requires $\mathcal{O}\!\left(l^{2}\right)$ time.
Combining the above yields a worst-case complexity of $\mathcal{O}\!\left(dkn+l\left(\min\!\left(2^{k},2^{l}\right)+\max\!\left(k,l\right)\right)\right)$. Since $k$ and $l$ are constant, the nearest-neighbor search dominates. The average total runtime of the search is $\mathcal{O}\!\left(dk\log n\right)$.
Appendix B Auxiliary Results
This section lists results underpinning the findings in Appendix A. Section B.1 introduces a closed-form expression of the expected probability mass. Section B.2 presents Lemma B.1, which is needed in the proof of Lemma 4.2.
B.1 Expected Probability Mass
This section, first, presents a closed-form expression for the expected probability mass of $\tilde{y}$ and $\tilde{y}_{\operatorname{c}}$. Second, we compare the computation with a simulation to numerically validate our closed-form expression and the results in Appendix A.
Closed-form expression. Let $\tilde{x}\in\operatorname{\mathcal{X}}$ be the instance of interest with hidden true label $\tilde{y}\in\operatorname{\mathcal{Y}}$ and $k$ associated neighbors $\left(x_{i},s_{i}\right)\in\operatorname{NN}_{k}\!\left(\tilde{x}\right)$. We recall that our label space $\operatorname{\mathcal{Y}}$ has $l\geq 3$ classes. Then, the expected probability mass of $\tilde{y}$ given instance $\tilde{x}$ is
$\ \operatorname{\mathbb{E}}_{\left(X,Y,S\right)\sim\operatorname{\mathbb{P}}}\!\big{[}\operatorname{\tilde{m}}\!\left(\left\{\tilde{y}\right\}\right)\mid X=\tilde{x}\big{]}\overset{(i)}{=}\operatorname{\mathbb{E}}_{\left(X,Y,S\right)\sim\operatorname{\mathbb{P}}}\!\big{[}\operatorname{q}\!\left(\left\{\tilde{y}\right\}\right)\mid X=\tilde{x}\big{]}$ (21)
=$\overset{(ii)}{=}$ $\ \operatorname{\mathbb{E}}_{\left(X,Y,S\right)\sim\operatorname{\mathbb{P}}}\!\Big{[}\sum_{\begin{subarray}{c}A_{1},\ldots,A_{k}\subseteq\operatorname{\mathcal{Y}}\\ \bigcap_{i=1}^{k}A_{i}=\left\{\tilde{y}\right\}\end{subarray}}\ \prod_{j=1}^{k}\operatorname{m}_{j}\!\left(A_{j}\right)\mid X=\tilde{x}\Big{]}$ (22)
=$\overset{(iii)}{=}$ $\ \underbrace{\sum_{h=0}^{k-1}\sum_{i=0}^{k-h-1}\sum_{j_{1}=0}^{k-h-1}\sum_{j_{2}=0}^{k-h-1}\dotsi\sum_{j_{l-2}=0}^{k-h-1}\!\!\!\!\!\!\!\sum_{\begin{subarray}{c}A_{1},\ldots,A_{k}\subseteq\operatorname{\mathcal{Y}}\text{;}\\ B_{1},\ldots,B_{l}\subseteq\left[k\right]\\ \text{with}\ |B_{1}|=h,|B_{2}|=i,\\ |B_{q+2}|=j_{q}\ \text{for}\ q\in\left[l-2\right]\text{;}\\ \tilde{y}\,\in\,A_{r}\ \text{for}\ r\,\in\,\left[k\right]\text{;}\\ A_{r}\,=\,\operatorname{\mathcal{Y}}\ \text{for}\ r\,\in\,B_{1}\text{;}\\ \tilde{y}_{\operatorname{c}}\in\,A_{r}\ \text{for}\ r\,\in\,B_{2}\text{;}\\ \tilde{y}_{q+2}\in\,A_{r_{q}}\ \text{for}\ r_{q}\,\in\,B_{q+2}\\ \text{and}\ q\,\in\,\left[l-2\right]\end{subarray}}}_{(iv)}\!\!\!\!\!\!\underbrace{\big{(}\prod_{j=1}^{k}\operatorname{m}_{j}\!\left(A_{j}\right)\big{)}}_{(v)}\,\underbrace{\big{(}\prod_{j=1}^{k}\operatorname{\mathbb{P}}\!\left(S=A_{j},Y=\tilde{y}\mid X=x_{j}\right)\big{)}}_{(vi)}$ (23)
$=$ $\ \sum_{h=0}^{k-1}\sum_{i=0}^{k-h-1}\sum_{j_{1}=0}^{k-h-1}\sum_{j_{2}=0}^{k-h-1}\dotsi\sum_{j_{l-2}=0}^{k-h-1}\!\!\!\!\!\!\!\underbrace{\sum_{\begin{subarray}{c}A_{1},\ldots,A_{k}\subseteq\operatorname{\mathcal{Y}}\text{;}\\ B_{1},\ldots,B_{l}\subseteq\left[k\right]\\ \text{with}\ |B_{1}|=h,|B_{2}|=i,\\ |B_{q+2}|=j_{q}\ \text{for}\ q\in\left[l-2\right]\text{;}\\ \tilde{y}\,\in\,A_{r}\ \text{for}\ r\,\in\,\left[k\right]\text{;}\\ A_{r}\,=\,\operatorname{\mathcal{Y}}\ \text{for}\ r\,\in\,B_{1}\text{;}\\ \tilde{y}_{\operatorname{c}}\in\,A_{r}\ \text{for}\ r\,\in\,B_{2}\text{;}\\ \tilde{y}_{q+2}\in\,A_{r_{q}}\ \text{for}\ r_{q}\,\in\,B_{q+2}\\ \text{and}\ q\,\in\,\left[l-2\right]\end{subarray}}}_{(vii)}\!\!\!\!\!\!\underbrace{\frac{1}{2^{k}}}_{(vi)}\underbrace{\left(\frac{1}{2^{l-2}-1}p_{1}\right)^{i}\left(\frac{1}{2^{l-2}}p_{2}\right)^{k-h-i}}_{(vi)}$ (24)
$=$ $\ \sum_{h=0}^{k-1}\sum_{i=0}^{k-h-1}\sum_{j_{1}=0}^{k-h-1}\sum_{j_{2}=0}^{k-h-1}\dotsi\sum_{j_{l-2}=0}^{k-h-1}\Bigg{[}\underbrace{\binom{k}{k-h}\binom{k-h}{i}\prod_{a=1}^{l-2}\binom{k-h}{j_{a}}}_{(vii)}$ (25)
$\quad-\underbrace{\binom{k}{k-h}\sum_{b=1}^{\min\left(i,j_{1},j_{2},\dots,j_{l-2}\right)}\left(-1\right)^{b+1}\binom{k-h}{b}\binom{k-h-b}{i-b}\prod_{c=1}^{l-2}\binom{k-h-b}{j_{c}-b}}_{(vii)}\Bigg{]}$ (26)
×$\ \times\frac{1}{2^{k}}\left(\frac{1}{2^{l-2}-1}p_{1}\right)^{i}\left(\frac{1}{2^{l-2}}p_{2}\right)^{k-h-i}\text{.}$ (27)
In the following, we show that transformations $(i)$ – $(viii)$ hold.
  • In $(i)$, we apply (4). For any $y\in\operatorname{\mathcal{Y}}$, it holds that $\operatorname{m}\!\left(\left\{y\right\}\right)=\operatorname{q}\!\left(\left\{y\right\}\right)$ as $\left\{y\right\}\neq\emptyset$ and $\left\{y\right\}\neq\operatorname{\mathcal{Y}}$.
  • In $(ii)$, we insert (4), that is, how all $\operatorname{m}_{i}$-s as defined by Algorithm 1 are combined into $\operatorname{q}$ using Yager’s rule (4).
  • In $(iii)$, we then compute the expected value by taking into account all possible combinations of sets $A_{1},\ldots,A_{k}\subseteq\operatorname{\mathcal{Y}}$ producing the intersection $\left\{\tilde{y}\right\}$.
  • Term $(iv)$ enumerates all cases where $\bigcap_{r=1}^{k}A_{r}=\left\{\tilde{y}\right\}$. To produce this intersection, label $\tilde{y}$ needs to be contained in all sets $A_{r}$. At most $k-1$ sets can be the full label space, that is, $h$ sets satisfy $A_{r}=\operatorname{\mathcal{Y}}$. The label $\tilde{y}_{\operatorname{c}}$ with which $\tilde{y}$ is most often confused has to be missing in at least one set $A_{r}$, that is, $i$ sets $A_{r}$ satisfy $\tilde{y}_{\operatorname{c}}\in A_{r}$. All remaining labels $\tilde{y}_{3},\ldots,\tilde{y}_{l}\in\operatorname{\mathcal{Y}}$ also have to be missing in at least one set $A_{r}$ each, that is, $j_{a}$ sets $A_{r}$ satisfy $\tilde{y}_{a+2}\in A_{r}$, respectively.
  • Term $(v)$ calculates the value that a given instantiation of all $A_{1},\ldots,A_{k}$ has. Because we only sum over focal elements of the $\operatorname{m}_{j}$-s ($j\in\left[k\right]$) by the construction of the sum in $(v)$, term $(vi)$ simplifies to $1/2^{k}$. Recall that all focal sets have a value of 1/2 (Algorithm 1, Line 6).
  • Expression $(vi)$ computes the probability that a given instantiation of all $A_{1},\ldots,A_{k}$ has. The variables $h$, $i$, $j_{1}$, …, $j_{l-2}$ directly allow to quantify the probabilities in a way to apply Assumption 4.1, which allows to simplify the expression.
  • As the summands within $(vii)$ do not depend on the $A_{1},\ldots,A_{k}$ anymore, we can simplify by counting their possible instantiations. We first count the possible combinations of sets where we pick to intersect $h$-times with the full label space and $\left(k-h\right)$-times with the neighbors’ candidate sets of which $i$ contain label $\tilde{y}_{\operatorname{c}}$ and $j_{a}$ contain the remaining labels for all $a\in\left[l-2\right]$. Second, we need to subtract all combinations that produce the whole label space within at least one candidate set as the probability of $s^{\prime}=\operatorname{\mathcal{Y}}$ is zero per Assumption 4.1. We do this using an inclusion-exclusion strategy. We have $k-h$ sets with which we intersect and that are not the full label space. Out of those $k-h$ sets, at least one set ought to contain all possible labels leaving $\binom{k-h-1}{i-1}$ and $\binom{k-h-1}{j_{c}-1}$ possibilities for the other sets. This overcounts combinations though as multiple sets can contain all labels. We employ the inclusion-exclusion strategy to remove combinations that produce the same allocations of labels to sets.
Similar to (27), we also express the expected mass of the most frequently co-occurring label $\tilde{y}_{\operatorname{c}}$ as
$\ \operatorname{\mathbb{E}}_{\left(X,Y,S\right)\sim\operatorname{\mathbb{P}}}\!\big{[}\operatorname{\tilde{m}}\!\left(\left\{\tilde{y}_{\operatorname{c}}\right\}\right)\mid X=\tilde{x}\big{]}=\operatorname{\mathbb{E}}_{\left(X,Y,S\right)\sim\operatorname{\mathbb{P}}}\!\big{[}\operatorname{q}\!\left(\left\{\tilde{y}_{\operatorname{c}}\right\}\right)\mid X=\tilde{x}\big{]}$ (28)
$=$ $\ \operatorname{\mathbb{E}}_{\left(X,Y,S\right)\sim\operatorname{\mathbb{P}}}\!\Big{[}\sum_{\begin{subarray}{c}A_{1},\ldots,A_{k}\subseteq\operatorname{\mathcal{Y}}\\ \bigcap_{i=1}^{k}A_{i}=\left\{\tilde{y}_{\operatorname{c}}\right\}\end{subarray}}\ \prod_{j=1}^{k}\operatorname{m}_{j}\!\left(A_{j}\right)\mid X=\tilde{x}\Big{]}$ (29)
$=$ $\ \sum_{h=0}^{k-1}\sum_{t=0}^{k-h-1}\sum_{j_{1}=0}^{k-h-1}\sum_{j_{2}=0}^{k-h-1}\dotsi\sum_{j_{l-2}=0}^{k-h-1}\Bigg{[}\binom{k}{k-h}\binom{k-h}{t}\prod_{a=1}^{l-2}\binom{k-h}{j_{a}}$ (30)
$\quad-{\binom{k}{k-h}\sum_{b=1}^{\min\left(t,j_{1},j_{2},\dots,j_{l-2}\right)}\left(-1\right)^{b+1}\binom{k-h}{b}\binom{k-h-b}{t-b}\prod_{c=1}^{l-2}\binom{k-h-b}{j_{c}-b}}\Bigg{]}$ (31)
×$\ \times{\frac{1}{2^{k}}\left(\frac{1}{2^{l-2}-1}p_{1}\right)^{t}\left(\frac{1}{2^{l-1}-1}p_{3}\right)^{k-h-t}}\text{.}$ (32)
To produce the intersection $\{\tilde{y}_{\operatorname{c}}\}$ in Yager’s rule, all $k-h$ sets with which we intersect need to contain $\tilde{y}_{\operatorname{c}}$. The variable $t$ denotes how many out of $k-h$ sets contain the true label $\tilde{y}$. Therefore, $\tilde{y}$ and $\tilde{y}_{\operatorname{c}}$ co-occur in a set in $t$ cases, with which the probability $(\frac{1}{2^{l-2}-1}p_{1})^{t}$ is associated. In $k-h-t$ sets, the true label is not contained, with which the probability $(\frac{1}{2^{l-1}-1}p_{3})^{k-h-t}$ is associated according to Assumption 4.1.
[…]
$\operatorname{\mathbb{P}}\!\left(S=\left\{1,2\right\},Y=1\mid X=x_{i}\right)$ =$\,p_{1}\text{,}$ (33)
$\operatorname{\mathbb{P}}\!\left(S=\left\{1\right\},Y=1\mid X=x_{i}\right)=\operatorname{\mathbb{P}}\!\left(S=\left\{1,3\right\},Y=1\mid X=x_{i}\right)$ $=\frac{1}{2}\,p_{2}\text{,}$ (34)
$\operatorname{\mathbb{P}}\!\left(S=\left\{2\right\},Y=2\mid X=x_{i}\right)=\operatorname{\mathbb{P}}\!\left(S=\left\{3\right\},Y=3\mid X=x_{i}\right)$ $=\frac{1}{3}\,p_{3}\text{,}$ (35)
$\operatorname{\mathbb{P}}\!\left(S=\left\{2,3\right\},Y=2\mid X=x_{i}\right)=\operatorname{\mathbb{P}}\!\left(S=\left\{2,3\right\},Y=3\mid X=x_{i}\right)$ $=\frac{1}{6}\,p_{3}\text{,}$ (36)
regarding a neighboring instance $x_{i}$ with $i\in\left[k\right]$. All other probabilities that are not listed are zero. Naturally, $\sum_{s^{\prime}\subseteq\operatorname{\mathcal{Y}}}\sum_{y^{\prime}\in\operatorname{\mathcal{Y}}}\operatorname{\mathbb{P}}\!\left(S=s^{\prime},Y=y^{\prime}\mid X=x_{i}\right)=p_{1}+2\frac{1}{2}p_{2}+2\frac{1}{3}p_{3}+2\frac{1}{6}p_{3}=1$.
Figure 2 compares the computation and simulation of the expected probability masses. The values of the computation and simulation are indiscernible. Note that the expected probability mass of any single-item set converges towards zero as $k$ increases. The more sets one intersects with in (4), the more likely it is to produce the empty set. As shown in Lemma 4.2 (part $(ii)$), the expected probability mass of $\tilde{y}$ is greater than the mass of $\tilde{y}_{\operatorname{c}}$, independently from the number of neighbors $k\in\mathbb{N}$.
B.2 Auxiliary Lemma to the Proof of Lemma 4.2
This section presents an auxiliary inequality, which is used in the proof of Lemma 4.2.
Lemma B.1.
Let $l\in\mathbb{N}$ with $l\geq 3$ denote the number of class labels and $k\in\mathbb{N}$ the number of neighbors. Let further $i,j_{a}\in\mathbb{N}_{0}$ with $0\leq i\leq j_{a}<k$ for all $a\in\left[l-2\right]$. Then, it holds that
$\sum_{b=1}^{i}\left(-1\right)^{b+1}\left(\frac{k}{k-b}\right)^{l-2}\binom{k}{b}\binom{k-b}{i-b}\prod_{c=1}^{l-2}\binom{k-b}{j_{c}-b}\ \geq\ \sum_{b=1}^{i}\left(-1\right)^{b+1}\binom{k}{b}\binom{k-b}{i-b}\prod_{c=1}^{l-2}\binom{k-b}{j_{c}-b}\text{.}$ (37)
Proof.
We first transform the statement as
$\ \sum_{b=1}^{i}\left(-1\right)^{b+1}\left(\frac{k}{k-b}\right)^{l-2}\binom{k}{b}\binom{k-b}{i-b}\prod_{c=1}^{l-2}\binom{k-b}{j_{c}-b}\ \geq\ \sum_{b=1}^{i}\left(-1\right)^{b+1}\binom{k}{b}\binom{k-b}{i-b}\prod_{c=1}^{l-2}\binom{k-b}{j_{c}-b}$ (38)
$\Leftrightarrow$ $\ \sum_{b=1}^{i}\left(-1\right)^{b+1}\left(\left(\frac{k}{k-b}\right)^{l-2}-1\right)\binom{k}{b}\binom{k-b}{i-b}\prod_{c=1}^{l-2}\binom{k-b}{j_{c}-b}\ \geq\ 0\text{.}$ (39)
We then write the latter as
$\ \sum_{b=1}^{i}\left(-1\right)^{b+1}\underbrace{\left(\left(\frac{k}{k-b}\right)^{l-2}-1\right)}_{(i)}\binom{k}{b}\binom{k-b}{i-b}\prod_{c=1}^{l-2}\binom{k-b}{j_{c}-b}$ (40)
$=$ $\ \sum_{b=1}^{i}\left(-1\right)^{b+1}\left(\sum_{c=0}^{l-3}\left(\frac{k}{k-b}\right)^{c}\right)\frac{b}{k-b}\binom{k}{b}\binom{k-b}{i-b}\underbrace{\prod_{c=1}^{l-2}\binom{k-b}{j_{c}-b}}_{(ii)}$ (41)
$=$ $\ \sum_{b=1}^{i}\left(-1\right)^{b+1}\left(\sum_{c=0}^{l-3}\left(\frac{k}{k-b}\right)^{c}\right)\underbrace{\frac{b}{k-b}\binom{k}{b}\binom{k-b}{i-b}\binom{k-b}{j_{1}-b}}_{(iii)}\prod_{c=2}^{l-2}\binom{k-b}{j_{c}-b}$ (42)
$=$ $\ \sum_{b=1}^{i}\left(-1\right)^{b+1}\left(\sum_{c=0}^{l-3}\left(\frac{k}{k-b}\right)^{c}\right)\frac{i}{k-j_{1}}\binom{k}{i}\binom{i-1}{b-1}\binom{k-b-1}{j_{1}-b}\prod_{c=2}^{l-2}\binom{k-b}{j_{c}-b}$ (43)
=$\overset{(iv)}{=}$ $\ \frac{i}{k-j_{1}}\binom{k}{i}\sum_{b=1}^{i}\left(-1\right)^{b+1}\left(\sum_{c=0}^{l-3}\left(\frac{k}{k-b}\right)^{c}\right)\binom{i-1}{b-1}\binom{k-b-1}{j_{1}-b}\prod_{c=2}^{l-2}\binom{k-b}{j_{c}-b}$ (44)
=$\overset{(v)}{=}$ $\ \frac{i}{k-j_{1}}\binom{k}{i}\sum_{b=0}^{i-1}\left(-1\right)^{b}\left(\sum_{c=0}^{l-3}\left(\frac{k}{k-b-1}\right)^{c}\right)\underbrace{\binom{i-1}{b}\binom{k-b-2}{j_{1}-b-1}}_{(vi)}\prod_{c=2}^{l-2}\binom{k-b-1}{j_{c}-b-1}$ (45)
$=$ $\ \frac{i}{k-j_{1}}\binom{k}{i}\sum_{b=0}^{i-1}\left(-1\right)^{b}\left(\sum_{c=0}^{l-3}\left(\frac{k}{k-b-1}\right)^{c}\right)\left(-1\right)^{-b}\binom{k-i-1}{j_{1}-1}\prod_{c=2}^{l-2}\binom{k-b-1}{j_{c}-b-1}$ (46)
=$\overset{(vii)}{=}$ $\ \frac{i}{k-j_{1}}\binom{k}{i}\binom{k-i-1}{j_{1}-1}\sum_{b=0}^{i-1}\left(\sum_{c=0}^{l-3}\left(\frac{k}{k-b-1}\right)^{c}\right)\prod_{c=2}^{l-2}\binom{k-b-1}{j_{c}-b-1}\geq 0\text{,}$ (47)
which shows the statement to be demonstrated as the last term is a sum of products whose factors are all non-negative. Recall that $0\leq i\leq j_{a}<k$ for all $a\in\left[l-2\right]$. In the following, we show that all transformations are sound.
  • Using Theorem 65, $(i)$ holds as
    $\left(\frac{k}{k-b}\right)^{l-2}-1$ $=\left(\left(\frac{k}{k-b}\right)^{l-2}-\left(\frac{k}{k-b}\right)^{l-3}\right)+\left(\left(\frac{k}{k-b}\right)^{l-3}-\left(\frac{k}{k-b}\right)^{l-4}\right)$ (48)
    $\qquad+\cdots+\left(\frac{k}{k-b}-1\right)$ (49)
    $=\left(\frac{k}{k-b}-1\right)\left(\left(\frac{k}{k-b}\right)^{l-3}+\left(\frac{k}{k-b}\right)^{l-4}+\cdots+1\right)$ (50)
    $=\left(\frac{k}{k-b}-1\right)\sum_{c=0}^{l-3}\left(\frac{k}{k-b}\right)^{c}=\frac{b}{k-b}\sum_{c=0}^{l-3}\left(\frac{k}{k-b}\right)^{c}\text{.}$ (51)
  • In $(ii)$, we extract the first factor of the product such that it starts with $c=2$.
  • Subsequently, we transform the binomial coefficients $(iii)$ as
    $\ \frac{b}{k-b}\binom{k}{b}\binom{k-b}{i-b}\binom{k-b}{j_{1}-b}$ (52)
    ${=}$ $\ \frac{b}{k-b}\binom{k}{i}\binom{i}{b}\binom{k-b}{j_{1}-b}$ (53)
    ${=}$ $\ \frac{b}{k-b}\frac{i!}{b!\left(i-b\right)!}\frac{\left(k-b\right)!}{\left(j_{1}-b\right)!\left(k-j_{1}\right)!}\binom{k}{i}$ (54)
    =$\overset{(viii)}{=}$ $\ \frac{i}{k-j_{1}}\frac{\left(i-1\right)!}{\left(b-1\right)!\left(i-b\right)!}\frac{\left(k-b-1\right)!}{\left(j_{1}-b\right)!\left(k-j_{1}-1\right)!}\binom{k}{i}$ (55)
    ${=}$ $\ \frac{i}{k-j_{1}}\binom{k}{i}\binom{i-1}{b-1}\binom{k-b-1}{j_{1}-b}\text{,}$ (56)
    where $(viii)$ rearranges terms.
  • Then, in $(iv)$, we move terms that are independent of $b$ outside of the summation.
  • In $(v)$, we shift the summation to start at $b=0$.
  • Finally, by defining $r=j_{1}-b-1\in\mathbb{N}_{0}$, we express $(vi)$ as
    $\binom{i-1}{b}\binom{k-b-2}{j_{1}-b-1}=$ $\ \binom{i-1}{b}\binom{r+k-j_{1}-1}{r}$ (57)
    ${=}$ $\ \left(-1\right)^{r}\binom{i-1}{b}\binom{j_{1}-k}{r}$ (58)
    =$\overset{(ix)}{=}$ $\ \left(-1\right)^{r}\sum_{\begin{subarray}{c}b^{\prime}\!,\,r^{\prime}\in\,\mathbb{N}_{0}\\ b^{\prime}+r^{\prime}=j_{1}-1\end{subarray}}\binom{i-1}{b^{\prime}}\binom{j_{1}-k}{r^{\prime}}$ (59)
    ${=}$ $\ \left(-1\right)^{j_{1}-b-1}\binom{i+j_{1}-k-1}{j_{1}-1}$ (60)
    ${=}$ $\ \left(-1\right)^{j_{1}-b-1}\left(-1\right)^{j_{1}-1}\binom{k-i-1}{j_{1}-1}$ (61)
    $=$ $\ \left(-1\right)^{-b}\binom{k-i-1}{j_{1}-1}\text{.}$ (62)
    Thereby, $(ix)$ holds as all summands that are introduced by $\sum_{b^{\prime}+r^{\prime}=j_{1}-1}\binom{i-1}{b^{\prime}}\binom{j_{1}-k}{r^{\prime}}$ are zero. The only non-zero summand is for $b^{\prime}=b$ and $r^{\prime}=r$, that is, $b+r=j_{1}-1$. In general, $\binom{m}{n}=0$ if $0\leq m<n$ for $m\in\mathbb{R}$ and $n\in\mathbb{N}_{0}$.
Figure 2: Comparison of simulating and calculating the expected belief / probability mass of the true label $\tilde{y}$ and of the most frequently co-occurring label $\tilde{y}_{\operatorname{c}}$ for $k\in\left[100\right]$, $l=3$, $p_{1}=0.4$, $p_{2}=0.35$, and $p_{3}=0.25$.
Appendix C External Theorems
This section collects identities and theorems that we use in the proofs of our statements.
Definition C.1.
(Binomial coefficients; Graham et al. 1994, p. 154). When both arguments of a binomial coefficient are natural, one defines
$\binom{n}{k}=\frac{n!}{k!\left(n-k\right)!}\text{,}$ (63)
for $n,k\in\mathbb{N}_{0}$ with $n\geq k\geq 0$. One extends this definition using falling factorials $r^{\underline{m}}:=\prod_{i=0}^{m-1}\left(r-i\right)$ for any $r\in\mathbb{R}$ and $m\in\mathbb{N}_{0}$. Binomial coefficients are then expressed as
$\binom{r}{m}=\frac{r^{\underline{m}}}{m!}\text{,}$ (64)
which allows for real arguments $r$.
Theorem C.2.
(Binomial theorem; Graham et al. 1994, p. 162). Given $x,y\in\mathbb{R}$ and $n\in\mathbb{N}_{0}$, the binomial theorem states
$\left(x+y\right)^{n}=\sum_{k=0}^{n}\binom{n}{k}x^{n-k}y^{k}\text{.}$ (65)
Theorem C.3.
(Graham et al. 1994, p. 164). As a special case of the binomial theorem, one obtains
$\left(-1\right)^{k}\binom{-r}{k}=\binom{r+k-1}{k}\text{,}$ (66)
for $r\in\mathbb{R}$ and $k\in\mathbb{N}_{0}$.
Theorem C.4.
(Graham et al. 1994, p. 168). Given $r\in\mathbb{R}$ and $m,k\in\mathbb{N}_{0}$ with $m\geq k$, it holds that
$\binom{r}{m}\binom{m}{k}=\binom{r}{k}\binom{r-k}{m-k}\text{.}$ (67)
Theorem C.5.
(Chu-Vandermonde identity; Graham et al. 1994, p. 169). Given $s,t\in\mathbb{R}$ and $n\in\mathbb{N}_{0}$, the Chu-Vandermonde identity states
$\binom{s+t}{n}=\sum_{k=0}^{n}\binom{s}{k}\binom{t}{n-k}\text{.}$ (68)
Appendix D Additional Experiments
This section contains additional experiments and their hyperparameters: Section D.1 provides experiments justifying Assumption 4.1, Section D.2 lists the values of all relevant hyperparameters, and the parameter sensitivity of our approach is discussed in Section D.3.
D.1 Applicability of Assumption 4.1
Given a fixed instance $\tilde{x}$, Assumption 4.1 describes the label distribution of the neighbors’ candidate sets. Thereby, we assume that the true label $\tilde{y}$ dominates the neighborhood. Label $\tilde{y}_{\operatorname{c}}$ co-occurs most often with the true label $\tilde{y}$. Otherwise, we assume uniform noise among the remaining noise labels.
Figure 3 demonstrates that those assumptions are satisfied for most of the four real-world datasets (see Section 5.3). In most cases, the true label $\tilde{y}$ occurs most often in the neighborhood of instance $\tilde{x}$. There are one or two other labels with which $\tilde{y}$ is commonly confused, which we model by $\tilde{y}_{\operatorname{c}}$. Apart from that, all remaining noise labels are distributed uniformly.
(a) Dataset bird-song.
(b) Dataset mir-flickr.
(c) Dataset msrc-v2.
(d) Dataset yahoo-news.
Figure 3: Counts of the class labels in the candidate sets $s_{i}$ of instance $\tilde{x}$’s 10-nearest neighbors $x_{i}$.
D.2 Hyperparameters
As mentioned in Section 5.1, we consider ten commonly used PLL approaches. We choose their parameters as recommended by the respective authors.
  • Pl-Knn (Hüllermeier & Beringer, 2005): For all non-MNIST datasets, we use $k=10$ neighbors as recommended by the authors. For the MNIST datasets, we use the hidden representation of a variational auto-encoder as instance features and use $k=20$. The variational auto-encoder has a 768-dimensional input layer (flat MNIST input), a 512-dimensional second layer, and 48-dimensional bottleneck layers for the mean and variance representations. The decoder uses a 48-dimensional first layer, a 512-dimensional second layer, and a 768-dimensional output layer with sigmoid activation. Otherwise, we use ReLU activations between all layers. Binary cross-entropy is used as a reconstruction loss. We choose the AdamW optimizer for training.
  • Pl-Svm (Nguyen & Caruana, 2008): We use the Pegasos optimizer (Shalev-Shwartz et al., 2007) and $\lambda=1$.
  • Ipal (Zhang & Yu, 2015): We use $k=10$ neighbors, $\alpha=0.95$, and 100 iterations.
  • Pl-Ecoc (Zhang et al., 2017): We use $L=\lceil 10\log_{2}(l)\rceil$ and $\tau=0.1$ as recommended.
  • Proden (Lv et al., 2020): For a fair comparison, we use the same base models for all neural-network-based approaches. We use a standard $d$-300-300-300-$l$ MLP (Werbos, 1974) for the non-MNIST datasets with ReLU activations, batch normalizations, and softmax output. For the MNIST datasets, we use the LeNet-5 architecture (LeCun et al., 1998). We choose the Adam optimizer for training.
  • Rc (Feng et al., 2020): We use the same base models as mentioned above for Proden.
  • Cc (Feng et al., 2020): We use the same base models as mentioned above for Proden.
  • Valen (Xu et al., 2021): We use the same base models as mentioned above for Proden.
  • Pop (Xu et al., 2023): We use the same base models as mentioned above for Proden. Also, we set $e_{0}=0.001$, $e_{end}=0.04$, and $e_{s}=0.001$. We abstain from using the data augmentations discussed in the paper for a fair comparison.
  • Dst-Pll (our proposed approach): Similar to Pl-Knn and Ipal, we use $k=10$ neighbors for the non-MNIST datasets. For the MNIST datasets, we use the hidden representation of a variational auto-encoder as instance features and use $k=20$. The architecture of the variational auto-encoder is the same as described above for Pl-Knn.
We have implemented all approaches in Python using the Pytorch library. All experiments need two to three days on a machine with 48 cores and one NVIDIA GeForce RTX 3090.
D.3 Parameter Sensitivity
Figure 4: Sensitivity of $k$ regarding the test-set MCC score, the fraction, and the MCC score of confident / non-rejected predictions.
Figure 4 shows the sensitivity of the number of neighbors $k$ regarding the test-set performance, the fraction of confident / non-rejected predictions, and the non-rejected prediction performance. The shaded areas indicate the standard deviation regarding the 5-fold cross-validation. As for default $k$-nearest neighbor classification, changes of $k$ have a relatively large impact. We show parameter sensitivity for each of the real-world datasets separately. Naturally, different datasets have different optimal parameter settings. The configuration $k=10$, which is also recommended within Pl-Knn (Hüllermeier & Beringer, 2005) and Ipal (Zhang & Yu, 2015), provides a good trade-off between the number of confident predictions and how accurate confident predictions are. Indeed, this setting produces a good number of confident predictions on most datasets (Figure 4; center plot). At the same time, it produces a good MCC performance of confident predictions on most datasets (Figure 4; right plot).
When increasing $k$, our method’s behavior on different datasets can be grouped into two groups. On the bird-song, mir-flickr, and yahoo-news datasets, increasing $k$ past ten neighbors also increases the amount of irrelevant labeling information from those neighbors. Therefore, our approach produces less confident predictions. At the same time, the MCC score of confident predictions remains at roughly the same level. This is because irrelevant labeling information from neighbors increases at most at the same rate as $k$. In contrast, on the lost and msrc-v2 datasets, the MCC score of confident predictions drops sharply at a certain point while the number of confident predictions decreases similarly. This is because irrelevant labeling information increases more rapidly than $k$: The decrease of confidence in predictions is slower than the increase of irrelevant candidate labels.